Operation Summary for September 26, 2011 - October 3, 2011 |
The Integrated Luminosity for the period from 9/26/11 to 10/3/11 was 49.66 pb-1 (inverse picobarns). NuMI reported receiving 8.21E18 protons on target during this same period. Friday September 30 The day shift began with Operations monitoring Tevatron (TEV) store 9158 (96.27E30) and the stash (291.6E10), with Pbar stacking (33.3E10), with the Main Injector (MI), Linac, and Booster operating normally, with the NuMI experiment off for an access to switch their horn polarity, with the MiniBooNE and Meson FTBF experiments T-1012 taking beam, and with the Collider Detector at Fermilab (CDF) and D-Zero (D0) using the colliding beams. During the day shift, operators began preparing Pbar for an access at 8:36 AM for media tours. Operations established beam to NuMI at 10:28 AM. At 2:36 PM, Helen Edwards terminated store 9158 and the TEV ramp; this marked the end of Collider Run II and the end of the Tevatron. Operators turned off everything else that needed to be turned off in the TEV. Operators prepared D0 and CDF for access. A RF person put the TEV RF system into standby. During the evening shift, Operations disabled beam to T-1012 at 6:55 PM. Pbar resumed stacking at 10:22 PM, but it tripped off at 11:58 PM due to a Debuncher kicker (D:IKIK) problem. Saturday October 1 The midnight shift began with Operations monitoring the stash (316.7E10) and the Pbar stack (28.2E10), and with NuMI and MiniBooNE taking beam. During the midnight shift, operators reported at 12:23 AM that D:IKIK problems continued; a Pbar person told operators to put Pbar into standby and a Kicker person would investigate the problem in the morning. At 2 AM, operators reported that all the Booster corrector magnet power supplies lost their ramps; operators restored a CAMAC crate. Operations established beam to T-1012 at 7:26 AM. During the day shift, a Pbar Kicker person investigated the problem with D:IKIK and changed out a high voltage power supply; Pbar resumed stacking at 11:49 AM. At 2:58 PM, a Switchyard power supply (V204) tripped off on a water temperature indication and took down beam to T-1012. They decided to end their experimental run at Fermilab. During the evening shift, operators halted stacking in Pbar at 7:32 PM; this marked the end of antiproton stacking at Fermilab. Sunday October 2 The midnight shift began with Operations monitoring the stash and stack, and with NuMI and MiniBooNE taking beam. During the midnight shift, at 5:44 AM, the vacuum went bad for Linac tank #1 and the I- Source vacuum valve closed; everything reset. Operators reported having problems with a Linac RF station (LRF5) all shift. During the day shift, operators prepared the TEV A-E sectors for access at 11:24 AM. During the evening shift, at 4:16 PM, ComEd called and said Fermilab had ten minutes before they were shutting down power to the Kautz Rd Substation due to a problem off site. Operators turned off MI. The stash was lost. The FESS Duty Electrician reported that the Kautz Rd power was back up by 5:41 PM. Operators began recovering vacuum and computer systems. Operations called in many people to help with recovery. Operators and a MI person brought the MI Dampers online. At 9:35 PM, Operators reported that they weren’t getting MDAT and TCLK signals from MI CAMAC crate #74; operators talked with a Controls person and ended up replacing a signal box. Operators reset the MiniBooNE BLM chassis and then established beam to MiniBooNE at 10:48 PM. Operators had to reset the NuMI BLM system also; NuMI resumed taking beam at 11:28 PM. Monday October 3 The midnight shift began with Operations monitoring the Pbar stack, and with the NuMI and MiniBooNE experiments taking beam. During the midnight shift, operators reset the MI BPM system at 12:09 AM. At 12:52 AM, a Booster RF station (BRF12) began sparking; operators bypassed the station and brought on BRF19 in its place. The MiniBooNE beam permit tripped off several times in an hour and ten minutes due to bad beam positions; it reset. Operators worked on recovering from the power outage all shift. Operations Coordinator (OPSCO) Plans Booster will conduct studies. Starting around 3 AM on Wednesday morning (10/5/11) NuMI will turn off in preparation for repositioning its target. |
|
More Information Comments and Suggestions |
| last modified 10/03/2011 email Fermilab |
| Security, Privacy, Legal | Use of Cookies | |